write down the maximum and minimum hexidecimal and decimal values achievable for each of the following data types in C:

uin8_t unsigned 8 bit integer

int8_t signed 8 bit integer

uin16_t unsigned 16 bit integer

int32_t signed 32 bit integer

explain everything
Answer #1

Unsigned 8-bit integer:

In case of 8-bit unsigned integer, we can represent integers from 0 to 28-1=255.

When all 8 bits are 0 then it represents 0 and when all 8 bits are 1 then it represents 255.

Hence minimum 8-bit unsigned integer: 0(decimal) and 0(hexadecimal)

Maximum 8-bit unsigned integer: 255(decimal ) and FF(hexadecimal).

Signed 8-bit integer:

In case of 8-bit signed number , most significant(left most) bit is used for sign bit(o for +ve and 1 for -ve).

So, positive numbers are 00000000 to 01111111. i.e. in decimal 0 to 27-1=127.

And negative numbers are, 10000000(-128) to 10000001(-1).

Hence minimum 8-bit signed integer: -128(decimal) and -80(hexadecimal)

Maximum 8-bit unsigned integer: 127(decimal ) and 7F(hexadecimal).

Unsigned 16-bit integer:

In case of 16-bit unsigned integer, we can represent integers from 0 to 216-1=65536.

When all 16 bits are 0 then it represents 0 and when all 16 bits are 1 then it represents 65536.

Hence minimum 8-bit unsigned integer: 0(decimal) and 0(hexadecimal)

Maximum 8-bit unsigned integer: 65536(decimal ) and 10000(hexadecimal).

Signed 16-bit integer:

In case of 16-bit signed number , most significant(left most) bit is used for sign bit(o for +ve and 1 for -ve).

So, positive numbers are 0000-0000-0000-0000 to 0111-1111-1111-1111. i.e. in decimal 0 to 215-1=32767.

And negative numbers are, 1000-0000-0000-0000(-32768) to 1000-0000-0000-0001(-1).

Hence minimum 16-bit signed integer: -32768(decimal) and -8000(hexadecimal)

Maximum 8-bit unsigned integer: 32767(decimal ) and 7FFF(hexadecimal).